### v4

# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversational Agents

Abstract
Recent advancements in Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents, making them applicable to various fields (e.g., education). Despite their progress, the evaluation of the agents often overlooks the complexities of real-world conversations, such as real-time interactions, multi-party dialogues, and extended contextual dependencies. To bridge this gap, we introduce DialSim, a real-time dialogue simulator. In this simulator, an agent is assigned the role of a character from popular TV shows, requiring it to respond to spontaneous questions using past dialogue information and to distinguish between known and unknown information. Key features of DialSim include assessing the agent’s ability to respond within a reasonable time limit, handling long-term multi-party dialogues, and evaluating performance under randomized questioning with LongDialQA, a novel, high-quality question-answering dataset. Our experiments using DialSim reveal the strengths and weaknesses of the latest conversational agents, offering valuable insights for future advancements in conversational AI. DialSim is available at https://dialsim.github.io/.
DialSim: A Real-Time Simulator for Evaluating Long-Term
Multi-Party Dialogue Understanding of Conversational Agents
Jiho Kim1, Woosog Chay1, Hyeonji Hwang1, Daeun Kyung1, Hyunseung Chung1, Eunbyeol Cho1, Yohan Jo2, Edward Choi1 1KAIST  2SNU {jiho.kim, edwardchoi}@kaist.ac.kr
1 Introduction
Figure 1: An overall process of DialSim. Gray speech bubbles indicate predetermined utterances from the script, and white speech bubbles indicate spontaneous questions asked during the simulation. Colored speech bubbles indicate the agent’s responses to the questions. (Left) An unanswerable question. (Center) A question that references a specific time. (Right) A multi-hop question that requires understanding past sessions (i.e., the Left and Center boxes). The question is asked in the format chosen by the user, either in a multiple-choice format or as an open-ended question.
Recent advancements in Natural Language Generation (NLG) within Large Language Models (LLMs) have significantly enhanced the capabilities of conversational agents. These agents are now integral to various fields, including entertainment (Zhou et al., 2023; Chen et al., 2024) and education (Ait Baha et al., 2023; Waisberg et al., 2024), providing personalized interactions that cater to individual preferences and interests. As they continue to evolve and become more widely adopted, it is crucial to rigorously assess their performance in real-world scenarios to ensure they meet user expectations and function effectively.
Traditionally, the evaluation of conversational agents has relied on qualitative assessments of their responses. This process typically involves human evaluators or LLMs judging the quality of an agent’s utterances (Adiwardana et al., 2020; Zhang et al., 2020; Roller et al., 2021; Shuster et al., 2022; Lee et al., 2023; Kim et al., 2024) or comparing responses between different agents on platforms like Chatbot Arena (Chiang et al., 2024). While these methods provide valuable insights into aspects such as naturalness and alignment with user instructions, they do not fully capture the complexities of real-world interactions.
In practice, conversational agents face a variety of challenges: engaging in real-time interactions, managing multi-party conversations, and recalling information from past dialogues. These scenarios demand more comprehensive evaluation methods—ones that test an agent’s ability to respond within a reasonable time constraint, understand multi-party dialogue contexts, and reason across extended interactions. To meet this demand, we introduce DialSim, a real-time dialogue simulator designed to evaluate the long-term multi-party dialogue understanding of conversational agents.
DialSim places the agent in the role of a main character within a TV show, engaging in extensive conversations based on the show’s scripted content (see Figure 1). During each session, a randomly selected character asks a randomly sampled question at an unpredictable time. The agent is evaluated on its ability to respond appropriately, relying solely on the dialogue history and acknowledging when it lacks sufficient information. This approach enables rigorous testing of dialogue comprehension in unpredictable, realistic scenarios. Additionally, the agent’s real-time interaction capabilities are assessed through time constraints for responses (e.g., 1s, 3s, 5s). To the best of our knowledge, this is the first work evaluating conversational agents under time constraints, introducing a novel dimension to agent performance assessment.
In order to run DialSim, a dialogue script and corresponding question-answer pairs are required. For this purpose, we created LongDialQA, a new question-answering dataset derived from long-term multi-party dialogues. It comprises dialogues from popular TV shows (i.e., Friends, The Big Bang Theory, and The Office), spanning approximately 1,300 sessions over five years, totaling around 350,000 tokens. Each session includes more than 1,000 questions curated through two approaches: refining questions from a fan quiz website and generating complex questions using extracted temporal knowledge graphs. ChatGPT-4 (OpenAI, 2023a) assisted in refining questions and extracting knowledge graphs, with all outputs meticulously reviewed to ensure quality.
LongDialQA also incorporates adversarial testing to rigorously challenge agents’ reliance on dialogue history rather than pre-trained knowledge. Since LLM-based agents may possess prior knowledge about the TV shows (see Appendix A), we developed adversarial tests that modify character names in two specific ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). These adversarial scenarios help verify that the agent’s responses are grounded in the contextual dialogue history rather than pre-trained knowledge.
Using DialSim, we evaluated the latest conversational agents, uncovering both their strengths and limitations. Our findings provide valuable insights for advancing conversational AI, emphasizing the need for robust, real-world evaluation frameworks.
2 Related Works
Conversational Agents Evaluation    Early evaluation methods for conversational agents often relied on reference-based metrics (e.g., BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)), which compare model outputs to gold dialogue references but often show weak correlation with human judgment Liu et al. (2016). In contrast, human evaluation—where human annotators assess coherence, factual correctness, consistency, and engagingness of the generated responses—provides reliable assessments Adiwardana et al. (2020); Zhang et al. (2020); Roller et al. (2021); Shuster et al. (2022); Lee et al. (2023), but it is costly and time-consuming.
With the advent of LLMs, new evaluation approaches have emerged. These include having LLMs evaluate utterances directly Li et al. (2023); Kim et al. (2024) or employing platforms (e.g., Chatbot Arena Chiang et al. (2024)) where humans rank responses from different agents. Despite these advances, existing methods are still limited to qualitative assessments of utterances and fail to capture real-world conversational scenarios (e.g., real-time interaction, and long-term multi-party dialogue). To address these limitations, we propose a dialogue simulator, DialSim, designed to evaluate a conversational agent’s comprehensive dialogue understanding capabilities.
Long-Term Dialogue Datasets    A representative dataset for long-term dialogue is Multi Session Chat (Xu et al., 2022), which features up to five sessions per dialogue. This dataset, created through crowdsourcing, ensures high-quality dialogues; however, generating longer dialogues via crowdsourcing has remained challenging. To address this issue, Conversation Chronicles (Jang et al., 2023) was developed by leveraging an LLM to create longer and more comprehensive conversational datasets. More recently, LoCoMo (Maharana et al., 2024) was created using both LLMs and crowdsourcing; it evaluates dialogue comprehension of an agent through various tasks (e.g., event summarization) in long-term dialogues. In contrast to other datasets generated through crowdsourcing or LLMs, LongDialQA leverages TV show scripts, naturally providing extended, multi-party dialogues that evolve over time. Building on these unique features, DialSim simulates realistic, long-term interactions to evaluate agents.
Datasets Based on the TV Show Scripts  While both TV show scripts and other dialogue datasets effectively capture dialogue characteristics, scripts offer a significant advantage due to their abundance and accessibility. This makes them particularly valuable for various dialogue understanding tasks such as question answering (QA) (Yang and Choi, 2019; Sang et al., 2022), coreference resolution (Chen and Choi, 2016; Chen et al., 2017; Zhou and Choi, 2018), relation extraction (Rashid and Blanco, 2018; Yu et al., 2020), and summarization (Gorinski and Lapata, 2015; Papalampidi et al., 2020; Chen et al., 2022). Notable datasets derived from scripts include FriendsQA (Yang and Choi, 2019) and TVShowGuess (Sang et al., 2022). FriendsQA treats each TV show scene as an independent conversation, with questions aiming to locate specific answer spans. TVShowGuess is a multiple-choice dataset requiring the identification of anonymized speakers in a scene based on prior context from earlier scenes. While many studies have utilized TV show scripts to create such datasets, only LongDialQA includes unanswerable questions and fully utilizes the extended context of scripts.
3 LongDialQA
To implement DialSim, we first developed LongDialQA, a question-answering dataset derived from long-term multi-party dialogues.
3.1 Data Construction
LongDialQA was developed using scripts from five consecutive seasons of popular TV shows (i.e., Friends, The Big Bang Theory, and The Office [^1] The scripts were downloaded from the website Kaggle (https://www.kaggle.com/).). These scripts were first preprocessed to serve as dialogue data (§ 3.1.1). Next, questions were generated for each script, drawing from fan quizzes (§ 3.1.2) and a temporal knowledge graph (TKG) (§ 3.1.3). Each question was then paired with the correct answer and multiple distractors. Finally, character style transfer was applied to refine the questions, resulting in the final pool of questions for each session (§ 3.1.4).
3.1.1 Script Preprocessing
The script we used includes 5 consecutive seasons per TV show, with each season containing approximately 20 episodes. Each episode is composed of multiple scenes (i.e., session). Each script includes not only utterances but also descriptions of characters’ actions and scenes, as well as metadata unrelated to the plot (e.g., names of writers and directors). We manually filtered out all irrelevant parts to create $Script_{pre}$, which contains only the conversations between characters. Additionally, since some of our questions involve time conditions (e.g., “Which friend wasn’t allowed to drive Monica’s Porsche in October 1994?”), we manually assigned a date to each scene in $Script_{pre}$ to provide time information to the agent. These dates were determined based on the contents of the conversations and the air dates of the episodes. The specific rules for date assignments are detailed in Appendix B. We then selected scenes involving the main character (i.e., Friends: Ross, The Big Bang Theory: Sheldon, The Office: Michael [^2] The characters with the most lines in each script were selected.) from $Script_{pre}$ and sequentially numbered them as sessions $S_{i}$. This process resulted in the final dialogue $\mathcal{D}=\{\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{N}\}$.
Figure 2: The overall process of question generation based on fan quizzes. First, we crawled fan quizzes from the web (1). Then, we applied filtering and revision processes to the crawled data (2-a, b). From this, we created secondary versions of the questions by adding dates to each (3-a). Then, we mapped each question to the scenes by determining whether it is answerable in that scene or not (3-b). Finally, we applied character style transfer to make the questions more natural (4).
3.1.2 Fan Quiz-Based Question Generation
We utilized a fan quiz website FunTrivia [^3] https://www.funtrivia.com/ to generate our questions. Fan quizzes cover a range of difficulty levels and focus on major events from each episode, making them promising for evaluating dialogue comprehension. Figure 2 illustrates our process for generating questions using fan quizzes. We began by extracting episode-specific quizzes from the site. Since these quizzes were created by dedicated fans, many required knowledge unrelated to the dialogue itself (e.g., “What is the name of the actor who played the clerk?”). To filter out these questions, we first selected quizzes that could be answered by referencing $Script_{pre}$ using ChatGPT-4 (OpenAI, 2023a). [^4] Fan quizzes exist for each episode, so we annotated them based on $Script_{pre}$ and then matched them to the sessions of $\mathcal{D}$. Questions about scenes without the main character are unanswerable, enabling us to design rigorous tests. Additionally, ChatGPT-4 annotated the scenes that served as evidence for each question. These annotations were verified by the authors to ensure accurate filtering and scene-mapping.
We then annotated the answerability of each question, i.e., whether it is possible for the main character to know the answer in the corresponding scene. For example, in Friends, if the evidence for a question was in scene 14, Ross would not know the answer if he was absent from that scene. Even if he were present in scene 14, he couldn’t answer the question if it had been asked in scene 1. However, if Ross appeared in scene 14 and the question was then asked in scene 15, he would know the answer. Using this principle, we determined whether each question is answerable. Additionally, to create questions that require long-term memory, new questions were generated by adding the date information of each scene to the questions (e.g., “How did Rachel buy her new boots on September 22, 1994?”). Detailed question generation processes are provided in Appendix C.
Figure 3: The overall process of question generation based on the temporal knowledge graph. We first extracted quadruples and constructed a temporal knowledge graph (1). Then, we generated questions based on this and mapped each question to the sessions by determining whether it was answerable in that session or not, similar to fan quiz-based questions (2-1, 2-2). Character style transfer was performed afterwards (3).
3.1.3 Temporal Knowledge Graph-Based Question Generation
Fan quizzes are useful for generating our questions, but since they are episode-specific and user-generated, the questions don’t span multiple episodes and their numbers are limited ($\sim$1K). To address this, we constructed a knowledge graph for each session and used it to generate questions. Initially, we used ChatGPT-4 to extract triples (i.e., [head, relation, tail]) from each session $S_{i}$ in $\mathcal{D}$. These triples were then refined by the authors. We employed 32 relations (e.g., girlfriend) derived from DialogRE (Yu et al., 2020), a high-quality dataset where human annotators manually extracted relations from Friends scripts, classifying relationships between characters into 37 categories. We adapted and modified these relations for our purpose. More details about the relations are provided in Appendix D.1. Finally, we combined the triples from each session with their respective dates to create a temporal knowledge graph (TKG) composed of quadruples (i.e., [head, relation, tail, date]).
Using the constructed TKG, we created questions that the main character could either answer or not for each session. We generated these questions by extracting one (i.e., one-hop) or two (i.e., two-hop) quadruples from the TKG. The form and answer of the question may change depending on the time it is asked, even if the same quadruple is used. For instance, if we select [Rachel, boyfriend, Ross, 1994-08-08] and ask the question in 1996, it would be: “Who was Rachel’s boyfriend on August 8th, 1994?” If asked on August 8th, 1994, the question would be: “Who is Rachel’s boyfriend?” In both cases, the answer is Ross. Conversely, if we inquire about Rachel’s boyfriend in 1992, when no information is available, the correct answer would be: “I don’t know.” In this manner, we manually verified the answer of each question. We applied the same principle to create more complex two-hop questions (e.g., “Rachel had a roommate on August 8th, 1994. Who is the boyfriend of the roommate now?”). The overall process of generating questions using TKG is illustrated in Figure 3. Examples of question templates and corresponding questions we created can be found in Appendix D.2.
3.1.4 Final Data Processing
Answer Choices Generation    To create multiple-choice questions, we carefully crafted a set of answer choices for each question. First, for all questions, we included a choice “(E) I don’t know.”, which agents must choose if the questions are unanswerable. For questions sourced from fan quizzes, the four answer choices were taken from the original quiz. The correct answers for these questions were the same as the original quiz, while the unanswerable questions were fixed to (E).
For TKG-based questions, the incorrect choices were derived from the tails of other quadruples that shared the same relation as the original quadruple. For example, for the question “Who is Rachel’s boyfriend?”, we extracted quadruples from the whole TKG where the relation is “boyfriend” and randomly selected three tails to form the incorrect choices. Additionally, to create a more adversarial test, if Rachel has a boyfriend in the past or future, we prioritized including these in the incorrect choices. In this case, for answerable questions (i.e., past or present), the correct answer is the tail of the original quadruple, while for unanswerable questions (i.e., future), the correct answer is (E).
  Friends The Big Bang Theory The Office
Total # of Tokens 335,439 367,636 352,914
Total # of Sessions 788 805 2,347
Fan Quiz Questions 192.9 26.7 42.7
TKG Questions 1173.2 1280.1 455.1
Question Candidates 1366.1 1306.8 497.9
[…]
Approx. # of Possible Tests $1366.1^{788}$ $1306.8^{805}$ $497.9^{2347}$
$\dagger$: Average number of questions per session
Table 1: Statistics of LongDialQA.
Question Style Transfer    In LongDialQA, questions are rephrased to reflect each character’s unique tone, creating the impression that the characters themselves are asking the questions (e.g., Generic style: “How did Rachel buy her new boots?$\rightarrow$ Style of Joey Tribbiani from Friends: “Hey, how did Rachel manage to snag those killer boots, huh?”). This transformation is powered by ChatGPT-4, and subsamples are reviewed by the authors to ensure that the original intent was preserved. More examples of style-transferred questions for each character are in Appendix E.
3.2 Statistics
Table 1 shows the statistics of LongDialQA.
4 DialSim
Building on LongDialQA, our simulator features an agent taking on the role of a main character in a dialogue (i.e., Ross, Sheldon, and Michael). Throughout the simulation, an agent is randomly asked questions by other characters that must be answered accurately within a time limit (§ 4.2).
4.1 Definition
Let the $k$-th utterance of the $n$-th session be denoted as $u_{n,k}$, and the $n$-th session consisting of $r$ utterances be $\mathcal{S}_{n}=\{\{u_{n,i}\}_{i=1}^{r},d_{n}\}$, where $d_{n}$ is the date of $\mathcal{S}_{n}$. The sub-session including up to the $k$-th utterance of the $n$-th session is $\mathcal{S}_{n,k}=\{\{u_{n,i}\}_{i=1}^{k},d_{n}\}$. The entire dialogue consisting of $N$ sessions is denoted as $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$. The agent’s memory up to the $k$-th utterance of the $n$-th session is $\mathcal{M}_{n,k}$. The agent answering question $q_{n,m,c}$ asked by character $c$ in the $m$-th utterance of the $n$-th session using the memory is $a_{n,m}=Agent(\mathcal{M}_{n,m},q_{n,m,c})$.
Input: $\mathcal{D}=\{\mathcal{S}_{i}\}_{i=1}^{N}$, Time interval $t$, Agent
Output: $C/T$ (CorrectAnswers / TotalQuestions)
1
[…]
3 $T$leftarrow 0 // TotalQuestions;
4 $\mathcal{M}_{1,0}\leftarrow\phi$;
5
6for $n\leftarrow 1$ to $N$ do
7       if $|Characters(\mathcal{S}_{n})|<2$ then
8             continue
9      else
10             $u_{n,m}\leftarrow SelectQuestionTiming(\mathcal{S}_{n})$;
11             $c\leftarrow RandCharInThreeTurns(u_{n,m})$;
12             $q_{n,m,c},a_{true}\leftarrow RandomQnA(n,m,c)$;
13             $T$leftarrow $T$ + 1;
14             for $k\leftarrow 1$ to $|\mathcal{S}_{n}|$ do
15                   $\mathcal{M}_{n,k}\leftarrow UpdateMemory(\mathcal{M}_{n,k-1},$ $u_{n,k},d_{n},t)$;
16                   if $k=m$ then
17                         $a_{n,m}\leftarrow AgentAnswer(\mathcal{M}_{n,m},$ $q_{n,m,c},d_{n},t)$;
18                         if $a_{n,m}=a_{true}$ then
19                               $C$leftarrow $C$ + 1;
20
21
22                  $\mathcal{M}_{n+1,0}\leftarrow\mathcal{M}_{n,k}$;
23
24
25
Algorithm 1 DialSim
4.2 Simulator
Algorithm 1 outlines the simulation process of DialSim, designed to emulate a real-time conversation. In this simulator, each participant’s utterance (including the agent’s) occurs at a predefined time interval (same as time limit), and the agent should update its memory within this interval. [^5] The memory can be incrementally updated in various ways (e.g., by storing each utterance separately or by summarizing the session up to the current utterance). A detailed discussion of these methods is provided in § 5.2. If updating the memory is not completed within the interval, the simulator will move on to the next utterance (Line 13). During the simulation, other characters ask questions (selected from LongDialQA) to the agent (Line 8-10), except in sessions where the agent is the only one talking (Line 5-6). The timing to ask a question is chosen randomly within the session (Line 8), and the speaker who asks the question is also chosen randomly. However, to make the simulation realistic, it is crucial to ensure that the chosen speaker is still present and hasn’t left the session. We achieved this by randomly choosing from characters who were present within three turns of the agent’s last utterance (Line 9). Then, a question is randomly selected and asked in the style of the corresponding speaker (Line 10). The agent then must respond to the question using its memory, all within the time limit (Line 15). The prompt for the response is created by combining the question with the dialogue history stored in the memory. If the response is not completed within the time limit, it will be considered a failure, and the simulator will move on to the next utterance. The prompt we used is provided in Appendix F.
5 Experiments
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 38.53 (0.89) 32.65 (2.65) 49.04 (1.67) 40.27 (1.36) 40.10 (0.75) 44.36 (2.36) 42.53 (1.26) ChatGPT-3.5 31.82 (1.31) 25.58 (1.78) 39.70 (1.86) 32.09 (0.84) 32.06 (1.60) 36.84 (1.77) 36.69 (1.25) Gemini 1.0 pro 2.96 (0.31) 28.77 (1.83) 25.07 (2.40) 35.27 (1.80) 34.22 (0.49) 31.83 (0.41) 35.75 (2.93) Open Tülu 2-70B 0.37 (0.15) 20.94 (0.75) 20.27 (0.99) 19.75 (0.08) 31.76 (1.84) 10.15 (0.55) 18.87 (0.30) Tülu 2-7B 0.84 (0.15) 12.68 (0.24) 19.58 (1.04) 26.84 (0.85) 14.08 (0.89) 17.39 (1.37) 25.21 (1.28) Llama3.1-70B 0.60 (0.06) 31.08 (1.21) 0.55 (0.12) 16.26 (5.05) 39.00 (0.30) 2.26 (0.42) 20.14 (0.22) Llama3.1-8B 28.82 (1.94) 27.12 (0.95) 34.14 (0.85) 30.91 (0.63) 29.76 (1.31) 33.25 (0.57) 24.48 (0.60) Mixtral-8x7B 1.88 (0.26) 16.84 (0.95) 26.23 (0.90) 17.11 (1.94) 17.94 (1.32) 26.78 (1.04) 15.40 (1.39) Mistral-7B 2.82 (0.46) 24.22 (2.04) 33.07 (1.01) 29.29 (1.76) 28.30 (1.93) 29.15 (1.67) 25.41 (1.53) Gemma-7B 16.60 (0.84) 22.11 (1.73) 24.30 (2.04) 18.33 (1.37) 26.42 (2.48) 22.54 (0.78) 18.80 (0.64) Gemma-2B 0.68 (0.20) 24.06 (2.03) 24.22 (1.34) 25.79 (1.00) 25.31 (1.55) 24.48 (1.62) 25.78 (1.12)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 2: The performance of the agents on Friends dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the highest performance for each retrieval method.
5.1 Experimental Setting
To efficiently and accurately evaluate the agents’ dialogue understanding abilities, we used a multiple-choice format for the questions in the experiments. Table 1 shows the statistics for LongDialQA, revealing a notable difference between the number of answerable and unanswerable questions. To ensure a balanced distribution of correct answers during the simulation, 20% of the questions were intentionally designed to be unanswerable, with each question offering five possible choices. In addition to the multiple-choice format, we also offer an option to use an open-ended format, allowing users to choose their preferred question format.
DialSim operates in real-time, requiring precise control of the experimental environment. Therefore, we conducted all experiments using the same hardware: NVIDIA RTX A6000 GPUs and an AMD EPYC 7702 64-Core Processor. The time limit used in the experiment was set to 6 seconds, based on the average time interval between utterances in the TV shows. Note that the time limit can be set to any value (even infinity) that meets one’s service requirement. We provide extensive discussions on the time limit feature of DialSim, including the test environment control and internet speed in Appendix G, along with details about question formats.
5.2 Baselines
We experimented with two methods for using an agent’s memory. The first method, namely Base LLM, is to simply prefix latest utterances as much allowed by the model’s context length. The second method, namely RAG-based, employs a retriever to search for relevant dialogue history from the agent’s memory (external storage) and includes it in the prompt (Lewis et al., 2020). This method can be broken down into three ways for storing dialogue history: each speaker’s utterance individually, the entire session, and a summarized version of each session (denoted as Utterance, Session Entire, and Session Sum. in Table 2). The retrieval from the memory was performed using BM25 (Robertson et al., 2009) and cosine similarity with the OpenAI embeddings (OpenAI, 2024c).
For the agents to be tested, we used both API-based models (i.e., Gemini-1.0 Pro, 1.5 Pro (Team et al., 2023; Reid et al., 2024), Claude 3 Opus (Anthropic, 2024), ChatGPT-3.5, 4o, 4o-mini (OpenAI, 2023b, 2024b, 2024a)) and open-source models (i.e., Tülu 2-7B, 70B (Ivison et al., 2023), Llama3.1-8B, 70B (Meta, 2024), Mistral-7B, 8x7B (Jiang et al., 2023, 2024), and Gemma-2B, 7B (Team et al., 2024)). [^6] Gemini-1.5 Pro, Claude 3 Opus, and ChatGPT-4o were evaluated only in the BM25-Session Entire and oracle setting to measure their performance upper bound due to their high prices. The experimental results can be found in Appendix K. To emulate conversational settings, we used chat templates for instruction-tuned models or directly used chat models.
5.3 Results
Overall Performance   Table 2 shows that API-based models outperformed open-source models due to their superior inference capabilities and faster response times in our setting. However, the performances of all baselines were below 50%, suggesting that current LLMs have limitations in their ability to serve as conversational agents for long-term multi-party dialogues. The experimental results for Friends, The Big Bang Theory, and The Office exhibited similar trends. The detailed results are described in Appendix H.
Model Base LLM BM25 OpenAI Embedding UT SE SS UT SE SS ChatGPT-4o-mini 38.91 34.44 49.21 42.23 38.91 43.64 42.40 ChatGPT-3.5 31.81 26.91 39.45 32.77 32.41 35.78 35.98 Gemini 1.0 pro 28.36 28.10 39.90 34.11 34.26 30.93 33.96 Llama3.1-70B 36.36 31.84 43.17 43.81 39.85 43.17 48.49 Llama3.1-8B 28.78 29.89 34.70 33.93 31.63 32.91 35.59 Mixtral-8x7B 42.19 31.84 46.47 32.31 35.51 41.24 34.18 Mistral-7B 32.93 28.20 35.09 30.16 30.12 31.00 30.10 $\dagger$: Limited the maximum context length to 8k tokens.
Table 3: The performance of the top-performing agents on Friends in DialSim without time limit. UT, SE, and SS denote Utterance, Session Entire, and Session Summary, respectively. Bold indicates the highest performance for each retrieval method, and underlined values indicate the best-performing retrieval method for each model. The full experimental results are in Appendix I.
For real-time interactions, selecting a model size that balances inference speed and reasoning ability is crucial. As shown in Table 2, under time constraints, differences in performance between model sizes often diminish, with smaller models sometimes outperforming larger ones due to faster inference. In contrast, as detailed in Table 3, larger models generally excel when no time limits are imposed, demonstrating superior reasoning capabilities. Interestingly, larger open-source models achieve inference performance comparable to API-based models, highlighting the trade-off between speed and accuracy. Therefore, selecting a model size that achieves a balanced trade-off is critical. Additional performance comparisons under varying time constraints are provided in Appendix J.
Storing the entire session consistently outperforms other history storing methods, as shown in Table 3. This is because individual utterances lack adequate context, and crucial information may be lost during summarization. However, Llama3.1 models achieved the best performance when using Session Sum. as a history saving method, owing to their strong summarization capabilities. Additionally, contrary to our expectations, Mixtral’s Base LLM (i.e., without history retrieval) outperforms some retrieval-based models in settings with unlimited time. This is due to Mixtral’s context length of 32k tokens, which is long enough to accommodate half a season of the script, allowing it to utilize a longer dialogue history than some of the other baselines. However, in a setting with a time limit, Mixtral’s performance significantly drops due to its long inference time. Therefore, for a conversational agent to converse in real-time, it is necessary to select a reasonably appropriate length of dialogue history.
Advanced techniques for storing and retrieving history are essential to engage in long-term multi-party dialogues. We conducted experiments under the oracle setting, where agents were given evidence sessions along with their dates (see Figure 2). Under these conditions, Llama3.1-70B achieved a top performance of 69.86% in an unlimited time scenario, outperforming the best RAG-based method by 21.37%. This significant performance gap highlights the importance of effective memory management techniques. Detailed experimental results are provided in Appendix K.
TKG-based questions present a greater challenge than fan quiz-based ones, with two-hop questions being particularly difficult. To assess the difficulty levels across different question types, we conducted an error analysis on ChatGPT-4o-mini, based on BM25-Session Entire, which showed the highest performance. The results showed that fan quiz-based questions had an accuracy of 58.80%, while TKG-based questions scored lower at 46.40%, highlighting the greater difficulty of TKG-based questions. Breaking down TKG-based questions further, one-hop questions had a performance of 66.67%, whereas two-hop questions had a performance of 13.53%, underscoring the challenge of two-hop questions. Furthermore, even in the oracle setting, while the performance of one-hop questions increased to 84.05%, two-hop questions remained at 28.45%. This indicates that two-hop questions are challenging not only in terms of history retrieval but also in reasoning across the given sessions.
Model Original Swapping Names New Names Llama3.1-70B 48.49 39.00 ($\downarrow$9.49) 44.27 ($\downarrow$4.22) Llama3.1-8B 35.59 31.59 ($\downarrow$4.00) 32.23 ($\downarrow$3.36) Mixtral-8x7B 46.47 37.72 ($\downarrow$8.75) 39.98 ($\downarrow$6.49) Mistral-7B 35.09 30.65 ($\downarrow$4.44) 34.59 ($\downarrow$0.50)
Table 4: The performance of the top-performing open-source agents on the adversarial test (without time limit). Numbers in parentheses represent the performance drop compared to the original test. The full experimental results are provided in Appendix L.
Adversarial testing is necessary to accurately evaluate dialogue understanding in conversational agents. We conducted further experiments for the adversarial test by altering the names of the characters in two ways: by swapping their names with each other (e.g., Joey $\leftrightarrow$ Monica) or by assigning new names to them (e.g., Joey $\rightarrow$ John). The results shown in Table 4 indicated a significant drop in overall performance compared to the original setup. This decline is attributed to the agents relying not only on the dialogue history but also on their pre-trained knowledge when answering questions. Additionally, the performance decrease was more pronounced when names were swapped compared to when new names were assigned. This suggests that new names represent new information, while mixed names in the dialogue history conflicted with the pre-trained knowledge, leading to reduced reasoning ability. The detailed experimental results are provided in Appendix L.
6 Conclusion
In this paper, we introduce DialSim, a simulator designed to evaluate the capabilities of conversational agents in understanding long-term, multi-party dialogues in real-time settings. To run DialSim, we first constructed LongDialQA, a dataset based on dialogues from well-known TV show scripts. LongDialQA also includes questions derived from fan quizzes and a temporal knowledge graph, enabling a comprehensive assessment of conversational agents. Using DialSim, we evaluated the latest conversational agents and uncovered significant limitations in their ability to effectively handle complex, multi-party, long-term dialogues in real-time scenarios.
Limitations
Despite its strengths, our simulator has two main limitations. First, while the questions and answers are logically paired for accurate evaluation, the random selection of questions could introduce a bit of awkwardness during conversations. Second, while we considered incorporating industry-specific dialogues such as chat logs from customer service or retail, where conversational agents could be used for business purposes, these dialogue datasets are usually proprietary and not publicly accessible. In future developments, we will focus on enhancing the natural flow of interactions and creating simulators that are applicable to real-world industries.
Appendix A LLM’s Prior Knowledge of the TV shows
We asked ChatGPT-4o to explain the plot of specific episodes of Friends. It accurately described the plots, as shown in Figure 45. Notably, it provided these answers without any web browsing, suggesting that ChatGPT-4o might have learned about these TV shows during its pre-training process.
Appendix B Date Assignment
We first extracted elements from the scripts that could indicate dates (e.g., Valentine’s Day, Christmas Eve). Then, we reviewed the scripts again to analyze the relative timing of the sessions. For example, if there is a line mentioning that Chandler broke up with his girlfriend two days ago, we annotated the session where he broke up with his girlfriend as occurring two days prior to the mentioned session. Next, while watching each episode, we pinpointed sessions where the dates might have changed by observing whether the characters’ outfits changed between sessions. Finally, we assigned a specific date to each session based on the actual broadcast date of the episode, adjusting for the relative differences in dates and events such as Christmas.
Appendix C Question Generation Based on Fan Quizzes
For each scene $s_{i,k}$ from episode $p_{i}$ in $Script_{pre}$, we define the set of answerable questions as $FanA_{i,k}$ and the set of unanswerable questions as $FanU_{i,k}$. The process of generating questions based on fan quizzes is as follows.
First, we collected quizzes for each season and episode of Friends, The Big Bang Theory, and The Office from the FunTrivia website. For each episode $p_{i}$ in $Script_{pre}$, we used ChatGPT-4 to determine if the crawled questions ${CrQ}_{i}=\{q_{i,0},q_{i,1},...,q_{i,l}\}$ could be answered using only $p_{i}$. If a question $q_{i,m}$ could be answered, ChatGPT-4 identified the scenes $ES_{i,m}$ that provide evidence for the answer, compiling them into $Q_{i}=\{(q_{i,m},ES_{i,m})\}_{m=0}^{l}$. Subsequently, the authors reviewed each $ES_{i,m}$, made necessary corrections, and annotated whether a single scene from $ES_{i,m}$ was sufficient to answer $q_{i,m}$ or if multiple scenes were needed to be considered simultaneously. For each $s_{i,k}$ within $p_{i}$, we assessed the answerability of the questions in $Q_{i}$.
For each $s_{i,k}$, if a question $q_{i,m}$ could be answered using just one scene, and $s_{i,k}$ occurs after the initial appearance of the main character in $ES_{i,m}$, we included $q_{i,m}$ in $FanA_{i,k}$. This ensures that the main character had adequate exposure to the relevant evidence. Additionally, for questions requiring verification across multiple scenes, if the main character appears in all $ES_{i,m}$ scenes and $s_{i,k}$ occurs after the last scene of $ES_{i,m}$, we included $q_{i,m}$ in $FanA_{i,k}$. If the main character does not appear in any of the $ES_{i,m}$ scenes, $q_{i,m}$ was included in $FanU_{i,k}$ since the main character has not experienced any evidence to answer the question. The rest are not included in the dataset as it is unclear whether they are answerable per scene. Additionally, to generate questions that require long-term memory, we added the most recent date of the evidence scenes for each question.
Appendix D Question Generation Based on a Temporal Knowledge Graph
D.1 Relations
We used the following 32 relations: ‘age’, ‘alumni’, ‘boss’, ‘boyfriend’, ‘brother’, ‘client’, ‘date of birth’, ‘dating with’, ‘ex-boyfriend’, ‘ex-fiance’, ‘ex-fiancee’, ‘ex-girlfriend’, ‘ex-husband’, ‘ex-roommate’, ‘ex-wife’, ‘father’, ‘fiance’, ‘fiancee’, ‘girlfriend’, ‘hometown’, ‘husband’, ‘job’, ‘major’, ‘mother’, ‘neighbor’, ‘pet’, ‘place of birth’, ‘place of work’, ‘roommate’, ‘sister’, ‘subordinate’, ‘wife’.
D.2 Question Templates and Generated Questions
Templates for one-hop questions are provided in Table 5 and Table 6. The former contains templates without temporal information, while the latter includes templates with temporal details. Since relations like “brother” and “sister” remain constant over time, questions about these relations do not require temporal information. Hence, no temporal templates were created for them. In Table 6, “on {time}” is used, but {time} can be not only the full date (year, month, and day) but also just the year and month, or even just the year. In these cases, “in {time}” is used.
The templates for two-hop questions are available in Table 7. These templates incorporate temporal information. To frame questions in the present tense, adjust the verbs to the present tense and remove the temporal information, following the approaches demonstrated in Table 5 and Table 6.
Appendix E Character Style Transfer
Table 8 shows the results of the character style transfer for three selected questions. To make the questions sound more natural and conversational, we prepended each one with “By the way,”. This helps them blend seamlessly into the flow of the conversation. The table shows how each question appears when rephrased in the style of various characters. The ‘Default’ setting is applied when the question is asked by a character who is not a recurring character of the TV show.
Appendix F Prompt for Response Generation
The prompt given to the conversational agent to answer questions using dialogue history is shown in Table 9. An example where the placeholders from Table 9 are filled with actual values can be found in Table 10.
Appendix G Experimental Setting
G.1 Time Limit
In DialSim, the time limit is a controllable parameter, giving developers the flexibility to conduct experiments with any chosen time constraint, or even without one. When a time limit is set, the experimental environment can impact performance. Consequently, depending on the environment in which the conversational agent is deployed, this could serve as a criterion for selecting the agent with relatively better performance. It is important to note that the primary objective of DialSim is not to evaluate the inference speed of LLMs, but rather to assess the end-to-end performance of conversational agents, where techniques like model sharding and tensor parallelism can be a part of the conversational agent to decrease the response latency if needed.
To control the environmental factors that could affect time, we conducted all experiments under the same conditions as described in Appendix G.1.1. The rationale for setting a 6-second time limit in our experiments is detailed in Appendix G.1.2, and an analysis of the Internet speed for API-based models can be found in Appendix G.1.3.
G.1.1 Environment Control
Our simulator operates in real-time, requiring precise control of the experimental environment. Therefore, we conducted all experiments using the same hardware: NVIDIA RTX A6000 GPUs and an AMD EPYC 7702 64-Core Processor. To maintain consistent CPU performance, we allocated 10 cores for each experiment and ensured that no other processes were running simultaneously.
G.1.2 Average Time Interval Between Utterances
Each episode includes around 240 utterances and lasts about 18 minutes without commercial breaks. This means each utterance should occur roughly every 4.5 seconds. However, because the experiments used the A6000, which is slower than the latest hardware like the A100 or H100, we extended the interval to 6 seconds.
To account for this, we set a 6-second window as the response time limit for agents and conducted experiments to determine whether current models could meet this criterion. It is important to emphasize that the primary goal of these experiments was not to evaluate the absolute performance of the models but to showcase the range of analyses possible under time limits.
G.1.3 Internet Speed
The performance of API-based models can be affected by internet speed. To analyze this, we conducted a comparative analysis of the response times between API-based models and open-source models. In our analysis of agents using OpenAI Embedding-Session Sum., we found that the API-based agents achieved average response times of 1.50 seconds for ChatGPT-4o-mini, 1.73 seconds for ChatGPT-3.5 and 2.69 seconds for Gemini 1.0 pro. In comparison, agents using open-source models showed average response times ranging from 2.06 seconds (Gemma 2B) to 7.15 seconds (Tulu2 70B). These results suggest that, even when accounting for both internet communication and model inference, remote API-based models are generally faster than open-source alternatives. This indicates that internet latency has a minimal impact on our evaluation.
G.2 Question Format
LongDialQA is a dataset that includes pairs of questions, answers, and choices. The questions are available in three formats: template-based multiple-choice, natural language multiple-choice, and open-ended. Users can choose any of these formats to evaluate the agent’s performance.
First, we provide multiple-choice questions in both template and natural language formats. For example, a template-based question might be, “Who was going out with Paul in September 1994?” with choices “(A) Emily, (B) Monica, (C) Ryan, (D) Rachel, (E) I don’t know”. In contrast, the same question in natural language format could be phrased as, “Who was going out with Paul in September 1994? Was it Emily, Monica, Ryan, Rachel, or do you not know?”
Additionally, we offer the option to ask questions in an open-ended format (e.g., “Who was going out with Paul in September 1994?”) without providing answer choices. This approach allows us to evaluate the agent’s ability to generate open-ended responses. The open-ended format is particularly useful for fan quiz-based questions, where some answers may require longer responses (e.g., Question: “Why did Monica and Chandler say they were late getting to the hospital?” Correct answer: “Monica went back for her jacket”).
For natural language multiple-choice and open-ended questions, a response is considered correct if it exactly matches the correct answer. If the response does not match exactly, the score is determined by comparing the response with the correct answer using a different language model (i.e., GPT-4o mini).
G.2.1 Choices in Multiple-Choice Questions
The number of questions based on fan quizzes was significantly smaller than the questions based on the TKG. Thus, 30% of the questions were intentionally extracted from the fan quiz-based during the simulation. Since each question has five choices, unanswerable questions were set to comprise 20% of the total to fairly stratify the correct answers.
G.3 Number of Retrieved Dialogue History
By default, agents retrieved up to 20 utterances, 10 entire sessions, and 15 session summaries, depending on the storing method, though some LLMs with shorter context lengths retrieved fewer histories accordingly.
Appendix H Experimental Results for The Big Bang Theory and The Office
The experimental results for The Big Bang Theory and The Office are provided in Table 11 and Table 12, respectively.
Appendix I Experimental Results in an Unlimited Time Setting
The experimental results for the unlimited time setting are presented in Table 13.
Appendix J Experimental Results for Different Time Limits
The experimental results for different time limits are shown in Figure 6 and Figure 7. Figure 6 illustrates the performance over different time limits in the BM25-Session Entire setting, while Figure 7 displays the performance in the Oracle setting. Due to the high costs, time-based experiments with ChatGPT-4o, Gemini-1.5 Pro, and Claude-3 Opus were conducted exclusively in the Oracle setting. One key observation from the results is the performance of ChatGPT-3.5, ChatGPT-4o-mini, and ChatGPT-4o. These models demonstrated consistent performance with quick inference times, handling up to a 3-second limit in the BM25-Session Entire setting and up to a 1-second limit in the Oracle setting. Consequently, these models are optimal for tasks requiring real-time communication without delays.
Appendix K Experimental Results in the Oracle Setting
Figure 8 shows the performance comparison between the BM25-Session Entire setting and the Oracle setting. These experiments were conducted without a time limit. Llama3.1-70B achieved the highest performance with a score of 69.86% in the Oracle setting.
Appendix L Experimental Results on Adversarial Test
In the adversarial test, we altered the characters’ names and ran experiments under different conditions. Table 14 displays the results when characters’ names were mixed with a 6-second time limit, while Table 15 shows the results without a time limit. Table 16 presents the results of changing characters’ names to new ones with a 6-second time limit, while Table 17 shows the results without a time limit.
Appendix M Annotator Instructions
Figure 9 and Figure 10 show the screenshots of the dataset labeling process. Figure 9 illustrates the annotation process for the questions based on fan quizzes, and Figure 10 describes the review process for selecting triples for the TKG.
Figure 4: The result of asking ChatGPT-4o to explain Season 2, Episode 7 of Friends.
Figure 5: The result of asking ChatGPT-4o to explain Season 3, Episode 14 of Friends.
Question Type Relation Template Question Example
[…]
Table 5: Templates for one-hop questions without temporal information.
Question Type Relation Template Question Example
[…]
Table 6: Templates for one-hop questions with temporal information.
First Relation Second Relation Template Question Example
roommate, wife, husband, girlfriend, boyfriend, client, neighbor, boss, subordinate, fiance, fiancee
roommate, wife, husband, pet,
girlfriend, boyfriend, client, neighbor,
boss, subordinate, fiance, fiancee
{sub1} had a {First Relation} on {time1}.
Who was the {Second Relation} of the
{First Relation} on {time2}?
Monica had a roommate on September 26th, 1994.
Who was the boyfriend of the roommate
on October 5th, 1996?
dating with
{sub1} had a {First Relation} on {time1}.
Who dated the {First Relation} on {time2}?
Monica had a roommate on September 26th, 1994.
Who dated the roommate on October 5th, 1996?
job, major, age
{sub1} had a {First Relation} on {time1}.
What was the {Second Relation} of the
{First Relation} on {time2}?
Monica had a roommate on September 26th, 1994.
What was the job of the roommate
on October 5th, 1996?
mother, father, son, daughter,
sister, brother
{sub1} had a {First Relation} on {time1}.
Who is the {Second Relation} of the
{First Relation}?
Monica had a roommate on September 26th, 1994.
Who is the mother of the roommate?
date of birth, place of birth,
{sub1} had a {First Relation} on {time1}.
When (Where) was the {First Relation} born?
Monica had a roommate on September 26th 1994.
When was the roommate born?
place of work
{sub1} had a {First Relation} on {time1}.
Where did the {First Relation} work
on {time2}?
Monica had a roommate on September 26th, 1994.
Where did the roommate work
on October 5th, 1996?
hometown
{sub1} had a {First Relation} on {time1}.
Where is the hometown of the
{First Relation}?
Monica had a roommate on September 26th, 1994.
Where is the hometown of the roommate?
dating with
roommate, wife, husband,
girlfriend, boyfriend, client, neighbor,
boss, subordinate, fiance, fiancee
{sub1} dated a person on {time1}.
Who was the {Second Relation} of the
person on {time2}?
Monica dated a person on September 26th, 1994.
Who was the boss of the person
on October 5th, 1996?
mother, father, son, daughter, sister, brother
roommate, wife, husband,
girlfriend, boyfriend, client, neighbor,
boss, subordinate, fiance, fiancee
Who was the {Second Relation} of {sub1}’s
{First Relation} on {time2}?
Who was the roommate of Ross’s
sister on September 26th, 1994?
dating with
Who dated {sub1}’s {First Relation}
on {time2}?
Who dated Ben’s father on September 26th, 1994?
job, age, major
What was the {Second Relation} of
{sub1}’s {First Relation} on {time2}?
What was the job of Ben’s father on September 26th, 1994?
mother, father, son, daughter,
sister, brother
Who is the {Second Relation} of
{sub1}’s {First Relation}?
Who is the mother of Ross’s son?
date of birth, place of birth
When (Where) was {sub1}’s {First Relation}
born?
When was Monica’s brother born?
place of work
Where did {sub1}’s {First Relation}
work on {time2}?
Where did Monica’s brother work
on October 5th, 1996?
hometown
Where is the hometown of
{sub1}’s {First Relation}?
Where is the hometown of Ross’s son?
Table 7: Templates for two-hop questions.
Original Question Character Style Transferred Question
By the way, how did Rachel buy her new boots? Default
Hey, any idea what Rachel used to snag those stylish
new boots of hers?
Monica
Hey, do you know what Rachel used to snag those super
cute new boots she’s been rocking?
Chandler
So, could we BE any more curious about how Rachel
snagged those new boots?
Joey
Hey, how did Rachel manage to snag those killer
boots, huh?
Phoebe
Oh my gosh! Do you have any idea how Rachel snagged
those super cute new boots?
By the way, who dated Monica on September 22, 1994? Default
So, who was Monica’s date on the night of
September 22, 1994?
Chandler
Oh, could you BE any more specific about who was
going out with Monica on September 22, 1994?
Joey
Hey, just outta curiosity, who was goin’ out with
Monica on September 22, 1994?
Phoebe
Oh my gosh, so like, who was Monica’s date on that
super specific day, September 22, 1994?
Rachel
Oh my god, so like, who was going out with Monica
on September 22, 1994?’
By the way, Rachel had a roommate on October 28, 1994. Who dated the roommate in September 1994? Default
Oh. My. God. Remember when Rachel had a roommate
back on October 28, 1994? So, who was going out with
that roommate by September 1994?
Monica
Hey, just out of curiosity, do you know who was going
out with Rachel’s roommate from back in September 1994?
I remember she got that roommate around October 28, 1994.
Chandler
So, just for a little stroll down memory lane, Rachel
was bunking with someone on October 28, 1994.
Any wild guesses on who was dating this mystery
co-habitant by September 1994?
Joey
Hey, so you know how Rachel was living with someone
back on October 28, 1994, right? So I’m just wonderin’
here, who was going out with this roommate of hers in
September 1994?
Phoebe
By the way, Rachel had a roommate on October 28, 1994.
Who dated the roommate in September 1994?
Table 8: Examples of the results of character style transfer.
Prompt for Response Generation
You are $<$Chatbot$>$, a long-term conversational agent capable of interacting with multiple users.
Based on the [Retrieved Dialogue History] provided, please answer the given [Question].
Note the following points:
1. Your answer must exclusively be one of the options: (A), (B), (C), (D), (E).
2. Your responses should solely rely on the retrieved dialogue history. If the information in the dialogue
history is insufficient to answer the question, you must choose (E).
3. This question is being asked in the context of $<$Date$>$.
[Retrieved Dialogue History]
$<$Dialog_History$>$
[Question] $<$Question$>$
[Answer]
Table 9: In the $<$Chatbot$>$ placeholder, the name of the main character (i.e., Ross, Sheldon, Michael) for each TV show is inserted. In the $<$Date$>$ placeholder, the date of the session in which the question is being asked is inserted. In the $<$Dialog_History$>$ placeholder, the dialogue history that the agent will use is inserted. In the $<$Question$>$ placeholder, the question that the agent should answer along with five choices is inserted.
Prompt for Response Generation
You are Ross, a long-term conversational agent capable of interacting with multiple users.
Based on the [Retrieved Dialogue History] provided, please answer the given [Question].
Note the following points:
1. Your answer must exclusively be one of the options: (A), (B), (C), (D), (E).
2. Your responses should solely rely on the retrieved dialogue history. If the information in the dialogue
history is insufficient to answer the question, you must choose (E).
3. This question is being asked in the context of [February 26, 1999].
[Retrieved Dialogue History]
[Session #1 on September 22, 1994]
<<Session Omitted>>
Ross: No, go on! It’s Paul the Wine Guy!
Phoebe: What does that mean? Does he sell it, drink it, or just complain a lot?
Monica: Hi, come in! Paul, this is.. … everybody, everybody, this is Paul.
All: Hey! Paul! Hi! The Wine Guy! Hey!
Chandler: I’m sorry, I didn’t catch your name. Paul, was it?
Monica: Okay, umm-umm, I’ll just–I’ll be right back, I just gotta go ah, go ah…
Ross: A wandering?
Monica: Change! Okay, sit down. Two seconds.
Phoebe: Ooh, I just pulled out four eyelashes. That can’t be good.
<<Session Omitted>>
[Session #2 on May 20, 1998]
<<Session Omitted>>
Rachel: Umm, hi!
Ross: Hi.
Rachel: Is Monica around? I-I have to ask her something.
Ross: She’s doing her laundry.
<<Session Omitted>>
Rachel: Y’know what Ross? You’re not going anywhere. You’re gonna sit right here.
I’m gonna make you a cup of tea and we’re gonna talk this thing whole out. All right? Hey, Dave!
Dave: Yeah?
Rachel: Umm, listen, I’m gonna need to take a rain check, my roommate is just really sick.
Okay? Bye! Honey, listen, I know, I know things seem so bad right now.
[Question] Chandler: So, just for a little stroll down memory lane, Rachel was bunking with someone in May 1998.
Any wild guesses on who was dating this mystery cohabitant by September 22, 1994?
(A) Paolo (B) Paul (C) Roger (D) Vince (E) I don’t know.
[Answer]
Table 10: An actual example of the prompt for response generation.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 22.68 (2.12) 19.77 (2.02) 36.63 (1.82) 30.10 (1.44) 29.54 (0.71) 32.34 (0.58) 35.72 (0.81) ChatGPT-3.5 32.49 (1.72) 25.32 (1.20) 35.59 (2.12) 33.86 (1.09) 27.81 (0.40) 32.97 (0.86) 37.02 (1.13) Gemini 1.0 pro 3.49 (0.69) 25.87 (1.23) 30.72 (0.18) 38.16 (1.25) 37.42 (0.68) 32.09 (0.44) 36.30 (0.32) Open Tülu 2-70B 0.62 (0.13) 21.08 (0.70) 18.95 (1.07) 22.36 (0.65) 34.64 (0.69) 9.08 (1.00) 20.22 (1.48) Tülu 2-7B 0.53 (0.18) 15.58 (1.34) 22.26 (0.53) 29.99 (0.57) 16.84 (2.13) 21.48 (0.77) 28.69 (1.15) Llama3.1-70B 0.25 (0.07) 21.55 (0.93) 0.15 (0.12) 1.26 (0.31) 34.21 (1.59) 3.89 (1.05) 14.89 (1.74) Llama3.1-8B 21.30 (1.68) 12.80 (1.06) 25.50 (0.16) 18.56 (0.99) 23.10 (2.69) 25.48 (3.65) 20.75 (1.58) Mixtral-8x7B 1.95 (0.34) 15.91 (0.71) 34.52 (1.12) 16.83 (1.60) 17.45 (0.49) 34.98 (0.99) 13.83 (2.18) Mistral-7B 3.11 (0.21) 24.69 (1.82) 34.26 (0.60) 32.17 (1.39) 30.23 (0.62) 33.36 (0.56) 29.19 (1.54) Gemma-7B 16.40 (0.74) 21.40 (2.33) 19.74 (2.45) 16.67 (0.40) 24.50 (1.87) 20.22 (1.39) 16.12 (0.52) Gemma-2B 1.56 (0.06) 28.94 (0.35) 26.12 (2.22) 33.47 (1.41) 27.92 (0.68) 29.40 (1.79) 34.86 (3.20)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 11: The performances of the agents on The Big Bang Theory dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 28.48 (1.01) 29.44 (0.62) 43.16 (1.37) 35.92 (2.50) 37.81 (0.30) 40.91 (0.37) 42.83 (1.12) ChatGPT-3.5 36.54 (0.32) 36.63 (0.57) 45.33 (1.00) 40.93 (0.13) 42.49 (1.24) 43.04 (0.82) 45.18 (0.56) Gemini 1.0 pro 2.42 (0.18) 35.11 (0.50) 48.90 (1.57) 40.91 (0.75) 44.72 (0.19) 46.63 (0.89) 45.82 (0.97) Open Tülu 2-70B 0.46 (0.09) 22.33 (1.00) 35.52 (0.89) 23.49 (1.16) 38.61 (1.02) 43.49 (1.27) 23.54 (0.52) Tülu 2-7B 0.32 (0.04) 25.86 (0.54) 27.95 (1.03) 36.60 (2.11) 22.13 (0.33) 29.50 (0.56) 35.51 (1.40) Llama3.1-70B 0.19 (0.07) 29.21 (0.56) 13.31 (0.94) 21.32 (5.22) 47.41 (0.93) 47.07 (1.32) 19.46 (1.73) Llama3.1-8B 21.87 (0.60) 22.03 (0.32) 37.94 (1.28) 29.16 (1.80) 27.76 (3.52) 37.67 (1.70) 26.67 (0.83) Mixtral-8x7B 1.53 (0.41) 19.63 (0.79) 34.35 (1.19) 16.07 (0.56) 20.02 (0.44) 30.44 (1.69) 12.43 (1.04) Mistral-7B 2.55 (0.09) 30.65 (0.45) 41.16 (1.26) 35.67 (1.68) 36.92 (2.13) 42.71 (1.24) 37.65 (2.42) Gemma-7B 17.81 (0.86) 21.58 (0.61) 25.62 (0.02) 12.20 (0.57) 24.88 (0.93) 24.38 (0.52) 15.70 (0.43) Gemma-2B 0.83 (0.16) 29.71 (0.69) 28.11 (1.14) 34.63 (0.94) 31.54 (0.65) 30.31 (0.16) 33.37 (0.27)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 12: The performances of the agents on The Office dialogue in DialSim (time limit = 6 seconds). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API ChatGPT-4o-mini 38.91 (0.99) 34.44 (0.52) 49.21 (0.12) 42.23 (1.57) 38.91 (0.74) 43.64 (0.42) 42.40 (0.99) ChatGPT-3.5 31.81 (1.33) 26.91 (2.30) 39.45 (1.40) 32.77 (1.31) 32.41 (0.96) 35.78 (0.74) 35.98 (1.75) Gemini 1.0 pro 28.36 (0.97) 28.10 (1.08) 39.90 (1.08) 34.11 (1.64) 34.26 (2.91) 30.93 (2.17) 33.96 (2.11) Open Tülu 2-70B 3.31 (0.32) 29.87 (0.65) 35.87 (2.56) 34.72 (1.63) 37.07 (0.72) 33.63 (1.32) 38.62 (1.94) Tülu 2-7B 1.57 (0.12) 28.93 (2.81) 28.72 (1.80) 30.86 (2.29) 34.55 (0.47) 31.04 (0.96) 32.12 (0.75) Llama3.1-70B 36.36 (0.68) 31.84 (1.29) 43.17 (0.99) 43.81 (0.94) 39.85 (2.08) 43.17 (0.68) 48.49 (0.97) Llama3.1-8B 28.78 (0.34) 29.89 (1.56) 34.70 (1.75) 33.93 (1.76) 31.63 (2.17) 32.91 (0.51) 35.59 (1.09) Mixtral-8x7B 42.19 (1.76) 31.84 (0.78) 46.47 (1.75) 32.31 (1.09) 35.51 (0.19) 41.24 (2.90) 34.18 (0.96) Mistral-7B 32.93 (0.59) 28.20 (1.17) 35.09 (1.76) 30.16 (1.82) 30.12 (1.45) 31.00 (1.93) 30.80 (1.75) Gemma-7B 18.78 (0.87) 22.26 (1.52) 23.62 (2.09) 19.83 (1.74) 25.07 (0.49) 22.48 (0.25) 20.08 (0.76) Gemma-2B 1.16 (0.26) 25.03 (1.54) 24.64 (1.31) 24.84 (2.05) 28.06 (1.38) 24.56 (2.60) 28.28 (1.94)
$\dagger$: Both ChatGPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 13: The performance of the agents on Friends dialogue in DialSim (without time limit). We conducted experiments three times and reported the accuracy and standard deviations.
Figure 6: The experimental results for different time limits in the BM25-Session Entire setting.
Figure 7: The experimental results for different time limits in the Oracle setting.
Figure 8: The performance comparison between the BM25-Session Entire setting and the Oracle setting.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 0.31 (0.13) 17.03 (0.94) 15.20 (0.87) 18.45 (1.04) 26.89 (0.54) 6.92 (0.47) 12.86 (1.23) Tülu 2-7B 0.73 (0.29) 12.50 (1.73) 17.58 (1.14) 24.21 (1.09) 10.20 (0.21) 14.26 (0.92) 21.03 (0.58) Llama3.1-70B 0.51 (0.00) 27.84 (1.89) 0.60 (0.06) 13.84 (2.31) 35.67 (1.89) 1.53 (028) 20.90 (0.37) Llama3.1-8B 25.84 (1.16) 25.24 (0.30) 28.86 (1.01) 24.56 (0.99) 28.86 (1.10) 32.35 (1.51) 24.05 (1.36) Mixtral-8x7B 1.77 (0.19) 14.03 (0.12) 21.11 (1.07) 15.50 (0.68) 13.14 (0.83) 18.03 (0.44) 18.47 (0.55) Mistral-7B 2.34 (0.17) 22.29 (1.43) 27.08 (0.99) 24.15 (1.76) 25.17 (1.74) 26.76 (2.64) 23.81 (2.53) Gemma-7B 18.87 (1.43) 22.85 (0.81) 22.96 (1.34) 17.95 (0.62) 25.46 (2.08) 21.53 (1.00) 17.66 (1.31) Gemma-2B 0.78 (0.22) 22.99 (0.66) 25.48 (1.54) 25.86 (2.48) 25.08 (1.34) 25.21 (0.22) 26.14 (1.71)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
Table 14: The performances of the agents on Friends dialogue in DialSim (time limit = 6 seconds, with shuffled names). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 2.54 (0.21) 26.47 (1.91) 31.75 (1.71) 30.94 (2.41) 31.90 (1.25) 29.83 (1.03) 31.86 (2.13) Tülu 2-7B 1.15 (0.06) 28.20 (1.63) 27.64 (2.37) 27.78 (1.32) 28.98 (0.96) 25.03 (0.94) 29.08 (2.47) Llama3.1-70B 31.38 (1.01) 29.08 (1.57) 36.48 (2.51) 36.91 (0.36) 35.89 (0.65) 39.80 (1.42) 39.00 (0.87) Llama3.1-8B 27.16 (1.62) 25.76 (1.42) 30.61 (1.25) 29.59 (1.25) 30.91 (0.99) 29.76 (1.26) 31.59 (0.69) Mixtral-8x7B 34.19 (0.68) 25.23 (1.19) 37.72 (0.96) 29.48 (0.87) 29.09 (1.46) 31.78 (1.71) 29.45 (0.04) Mistral-7B 27.78 (1.62) 25.02 (1.26) 30.65 (1.39) 24.99 (1.51) 27.34 (0.49) 27.97 (1.31) 26.97 (1.45) Gemma-7B 17.98 (2.15) 21.64 (0.39) 22.31 (2.15) 18.66 (1.55) 25.97 (1.92) 21.79 (0.40) 21.22 (0.59) Gemma-2B 1.04 (0.19) 24.19 (0.82) 25.25 (1.02) 24.32 (1.55) 25.03 (0.66) 25.44 (1.96) 23.62 (0.36)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
Table 15: The performances of the agents on Friends dialogue in DialSim (without a time limit and with shuffled names). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 0.21 (0.07) 18.24 (0.84) 20.60 (1.00) 18.64 (1.81) 31.71 (2.22) 7.82 (1.57) 17.31 (0.61) Tülu 2-7B 0.74 (0.15) 13.19 (0.35) 19.54 (1.29) 26.07 (2.00) 13.87 (0.71) 18.35 (1.21) 27.48 (2.04) Llama3.1-70B 0.64 (0.10) 29.29 (0.59) 0.60 (0.12) 15.07 (5.12) 39.08 (0.99) 2.43 (0.10) 18.18 (0.16) Llama3.1-8B 26.61 (1.24) 26.86 (0.78) 31.20 (1.87) 27.08 (0.63) 24.82 (1.07) 31.72 (1.66) 22.69 (0.63) Mixtral-8x7B 2.41 (0.40) 14.90 (0.82) 23.55 (0.40) 15.64 (0.47) 16.43 (1.68) 22.95 (0.68) 13.22 (1.61) Mistral-7B 3.35 (0.58) 24.44 (1.13) 31.39 (0.70) 24.26 (1.60) 29.82 (0.95) 30.21 (0.90) 23.90 (0.51) Gemma-7B 18.05 (0.97) 22.52 (0.81) 20.64 (0.26) 16.63 (1.59) 23.41 (1.26) 18.34 (0.82) 19.48 (2.45) Gemma-2B 0.47 (0.13) 24.31 (0.96) 24.77 (0.74) 25.74 (1.46) 28.41 (1.20) 24.68 (1.45) 24.75 (1.50)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
Table 16: The performances of the agents on Friends dialogue in DialSim (time limit = 6 seconds, with new names replaced). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. Open Tülu 2-70B 2.17 (0.46) 27.24 (1.17) 33.34 (1.17) 32.85 (1.85) 34.95 (0.47) 29.41 (1.22) 33.55 (2.79) Tülu 2-7B 0.63 (0.26) 30.26 (1.03) 27.68 (1.24) 30.98 (1.08) 30.99 (0.22) 27.93 (1.97) 31.80 (2.05) Llama3.1-70B 31.03 (1.91) 28.91 (2.33) 38.44 (5.98) 41.68 (3.68) 38.40 (1.10) 40.83 (1.07) 44.27 (0.57) Llama3.1-8B 26.65 (1.19) 25.80 (0.18) 32.01 (1.10) 30.48 (1.30) 29.50 (1.10) 32.82 (0.63) 32.23 (2.32) Mixtral-8x7B 38.92 (1.61) 26.91 (1.46) 39.98 (2.98) 27.02 (0.42) 30.27 (1.37) 35.89 (0.37) 29.52 (1.28) Mistral-7B 29.10 (1.34) 23.33 (0.83) 34.59 (0.80) 27.87 (2.66) 30.59 (2.09) 30.45 (0.89) 27.93 (0.99) Gemma-7B 17.37 (0.77) 22.58 (1.62) 21.41 (1.53) 21.61 (1.53) 23.90 (0.90) 21.61 (1.09) 20.88 (0.91) Gemma-2B 0.37 (0.07) 23.48 (1.85) 23.13 (1.14) 25.72 (2.67) 29.12 (1.90) 24.88 (1.40) 24.81 (1.31)
$\dagger$: Llama3.1 supports up to 128k tokens, but we limited it to 8k tokens due to GPU VRAM limits.
Table 17: The performances of the agents on Friends dialogue in DialSim (without a time limit and with new names replaced). We conducted experiments three times and reported the accuracies and the standard deviations.
Figure 9: The actual process of annotating questions from fan quizzes.
Figure 10: The actual process of reviewing extracted triples.